### v2

@@ 1 Introduction
    deterministic or (variants of) probabilistic computation
@@
    null or positive (measure) domain or range.
@@
Under randomness-preservation (in the sense of Martin-Löf [20]) injections are easier to invert. However the existence of an injective partial computable oneway function is unknown. A partial answer is given in §4 where a left-oneway (the notion of Gács [11]) partial computable injection is constructed.
Notation. We use $x,y,\cdots$ to denote reals and $\sigma,\tau,\cdots$ to denote strings. The bit of $x$ at position $n$ is $x(n)$, starting with $n=0$. Let $|\sigma|$ be the length of string $\sigma$. Let $2^{\omega}$ be the collection of all reals and $2^{<\omega}$ the collection of all strings. Let $\preceq$ be the prefix relation between strings and reals, and $\prec$ the strict prefix relation. Let $0^{n}$ be the $n$-bit string of 0s and $0^{\omega}:=000\dots$. For $x,y\in 2^{\omega}$ let
  • $x\upharpoonright_{n}$ be the $n$-bit prefix of $x$ so $x\upharpoonright_{n}=x(0)x(1)\cdots x(n-1)$
  • $x\oplus y$ be the real $z$ with $z(2n)=x(n)$ and $z(2n+1)=y(n)$
A tree $T$ is a non-empty subset of $2^{<\omega}$ such that if $\sigma\prec\tau\in T$ then $\sigma\in T$. A real $x$ is a path of $T$ if $\forall n,\ x\upharpoonright_{n}\in T$ Let $[T]$ be the class of all paths of $T$.
The Cantor space is $2^{\omega}$ with the topology generated by basic open sets
$\llbracket\sigma\rrbracket=\{x\in 2^{\omega}:\sigma\prec x\}$
for $\sigma\in 2^{<\omega}$ which are also clopen. Let $\mu$ be the uniform measure on $2^{\omega}$ defined by $\mu(\llbracket\sigma\rrbracket)=2^{-|\sigma|}$. We identify $2^{\omega}$ with $2^{\omega}\times 2^{\omega}$ via $(x,y)\mapsto x\oplus y$. Classes of reals with positive $\mu$-measure are positive and those with $\mu$-measure 00 are null.
@@
We use the notion of randomness and relative randomness in the sense of Martin-Löf [20] and denote the class of all random reals by $\mathsf{ML}$. We often refer to van Lambalgen’s Theorem [27] (see [8, Corollary 6.9.3]): for everyl $x,y,w$ the real $x\oplus y$ is $w$-random if and only if $x$ is $w$-random and $y$ is $x\oplus w$-random.
@@ 2 Complexity bounds and inversions
We give complexity bounds on the domain, range and inversions of partial computable functions, assuming familiarity with the arithmetical, hyperarithmetical and analytic hierarchies on $\mathbb{N},2^{\omega}$. We write $\Pi^{0}_{1}(w)$ for $\Pi^{0}_{1}$ relative to oracle $w$ and similarly for the other definability classes.
@@ 2.1 Partial computable functions on the reals
A Turing machine with an infinite oracle tape and an infinite one-way output tape can be regarded as an effective map $f$ on the reals: given $x\in 2^{\omega}$ on its oracle tape the machine runs indefinitely and if the cells of the output tape contain the digits of $y$ we write $y=f(x)$. This is the standard concept of computability on the reals from computable analysis [23, 28]. Let
  • $f:\subseteq 2^{\omega}\to 2^{\omega}$ denote a function from a nonempty subset of $2^{\omega}$ to $2^{\omega}$.
  • $\mathsf{dom}(f)$ be the domain of $f$ and $f(2^{\omega})$ be the range of $f$.
@@
A representation is a partial computable $\hat{f}:\subseteq 2^{<\omega}\to 2^{<\omega}$ with
@@
and is a representation of $f:\subseteq 2^{\omega}\to 2^{\omega}$ if $\hat{f}(x\upharpoonright_{n})\prec f(x)$ for all $x,n$ and
$f(x)\downarrow\iff\lim_{n}|\hat{f}(x\upharpoonright_{n})|=\infty.$
A function $f$ is $w$-computable if and only if it has a $w$-computable representation.
@@ 2.2 Complexity of domain and range
If $f:\subseteq 2^{\omega}\to 2^{\omega}$ is partial computable:
@@
$P=\bigcap_{i}\bigcup_{j}V_{i,j}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptV_{0,0}=2^{\omega},V_{i,j}\subseteq V_{i,j+1}.$
@@
Continuous functions $f$ on subsets of $2^{\omega}$ map compact sets to compact sets. Since $2^{\omega}$ is a Hausdorff space, the same holds for closed sets. An effective version of this fact is true when $f$ is partial computable.
@@
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be partial computable. For any $w\in 2^{\omega}$:
@@
For (i) let $\hat{P}$ be a $w$-computable tree with $[\hat{P}]=P$. We show
$y\in f(P)\iff\forall n\ \exists\sigma\in\hat{P}\cap 2^{n},\ \hat{f}(\sigma)\prec y.$ (1)
If $y\in f(P)$ then $\exists x\in P,\ f(x)=y$ so $\forall n\ \big{(}{x\upharpoonright_{n}\in\hat{P}\cap 2^{n}\ \wedge\ \hat{f}(x\upharpoonright_{n})\prec y}\big{)}$.
For the converse, suppose that there is $\sigma_{0},\sigma_{1},\dots$ with $\sigma_{i}\in\hat{P}\cap 2^{i}$ and $\hat{f}(\sigma_{i})\prec y$. Since $2^{\omega}$ is compact, there is a subsequence of $\sigma_{i}$ (without loss of generality, the sequence $\sigma_{i}$ itself) that converges to some $x$. Let $\tau_{i}$ be the longest common prefix of $\sigma_{i}$, $x$. Since $\sigma_{i}\to x$, we get $|\tau_{i}|\to\infty$ so $\tau_{i}\to x$. Also
  • since $\sigma_{i}\in\hat{P}$ we get $\tau_{i}\in\hat{P}$ so $x\in[\hat{P}]=P\subseteq\mathsf{dom}(f)$ and $f(x)\downarrow$
@@
Although the domain of any partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ is arithmetically definable (indeed $\Pi^{0}_{2}$) as a subset of $2^{\omega}$ it is possible that it is nonempty and has no arithmetically definable member.
Proposition 2.3.
The following hold for partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$:
  1. (i)
    if $x\in\Delta^{1}_{1}$ there is $f$ with domain $\mathsf{dom}(f)=\{z\}$ and $x<_{T}z$
  2. (ii)
    there is $f$ with uncountable $\mathsf{dom}(f)$ and $\mathsf{dom}(f)\cap\Delta^{1}_{1}=\emptyset$.
Proof..
For each $x\in\Delta^{1}_{1}$ there is an ordinal $\alpha<\omega_{1}^{\rm CK}$ such that $x<_{T}\emptyset^{(\alpha)}$. By [7, Theorem 2.1.4] $\{\emptyset^{(\alpha)}\}$ is $\Pi^{0}_{2}$. So (i) holds by Proposition 2.1 (iii). Let
be a computable injection from $\omega^{\omega}$ into $2^{\omega}$. By [7, Proposition 2.5.2] there is a computable tree $T\subseteq\omega^{\omega}$ such that $[T]$ is uncountable and $[T]\cap\Delta^{1}_{1}=\emptyset$.
Every $x\in\pi([T])$ computes a member of $[T]$ so $\pi([T])\cap\Delta^{1}_{1}=\emptyset$. Since the image $\pi([T])$ is a $\Pi^{0}_{2}$ subset of $2^{\omega}$ we get (ii) by Proposition 2.1 (iii). ∎
In particular a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ may have a non-empty domain such that all arithmetically definable closed subsets (represented as the set of infinite paths through an arithmetically definable tree) of it are empty.
By Proposition 2.2 (i) the range of any total computable $f:2^{\omega}\to 2^{\omega}$ is $\Pi^{0}_{1}$. On the other hand the range of a partial computable $f$ is only $\Sigma^{1}_{1}$. The latter follows by the representation of $\Sigma^{1}_{1}$ classes with quantifiers over the Baire space $\omega^{\omega}$ and an effective embedding of $\omega^{\omega}$ into $2^{\omega}$.
@@
    the range of any partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ is $\Sigma^{1}_{1}$
@@
    each $\Sigma^{1}_{1}$ class is the range of a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$.
@@
Clause (i) is straightforward. For (ii) let $Q\in\Sigma^{1}_{1}$.
There exists a total Turing functional $\Phi$ such that
  • $\forall\alpha\in\omega^{\omega}$ there exists at most one $y\in 2^{\omega}$ with $\forall s,\ \Phi(\alpha\oplus y)(s)=1$
  • $y\in Q\iff\exists\alpha\in\omega^{\omega}\ \ \forall s:\ \Phi(\alpha\oplus y)(s)=1$.
Since $\Phi$ is total the function $\varphi:\subseteq\omega^{\omega}\to 2^{\omega}$ given by
$\varphi(\alpha)=y\iff\forall s:\ \Phi(\alpha\oplus y)(s)=1$
is partial computable. Define an effective injection from $\omega^{\omega}$ to $2^{\omega}$ by
@@
Then $f(x):=\varphi(\pi^{-1}(x))$ is partial computable and $f(2^{\omega})=Q$. ∎
@@
If $f:\subseteq 2^{\omega}\to 2^{\omega}$ is partial computable with positive domain
@@ 2.3 Complexity of inversion
If $f,g:\subseteq 2^{\omega}\to 2^{\omega}$ we say $g$ is an inversion of $f$ if $\forall y\in f(2^{\omega}),\ f(g(y))=y$.
By [2, Corollary 3.2] every total computable injection $f$ has a total computable inversion. We are particularly interested in functions that preserve randomness.
@@
We say that $f:\subseteq 2^{\omega}\to 2^{\omega}$ is random-preserving if it has positive domain and $f(x)$ is random for each random $x\in\mathsf{dom}(f)$.
@@
$p^{z}(\langle n,s\rangle)$ :={ $:=\ \begin{cases}\ \ 2n&\textrm{if $n\in z^{\prime}_{s+1}-z^{\prime}_{s}$}\\[5.69046pt] 2\langle n,s\rangle+1&\text{otherwise}\end{cases}$
$h^{z}(x)(\langle n,s\rangle)$ $:=\ x(p^{z}(\langle n,s\rangle))$
@@
is a total computable random-preserving surjection. For contradiction, assume that $g$ is a continuous inversion of $f$. Without loss of generality we suppose that $\mathsf{dom}(g)=f(2^{\omega})$ which is $\Pi^{0}_{2}$, so $g$ has $G_{\delta}$ domain and therefore must be $w$-computable for some $w$. Now
@@
Given $f,g:\subseteq 2^{\omega}\to 2^{\omega}$ we say that $g$ is a
@@
By [2, Theorem 4.5] there is a total computable random-preserving surjection $f$ such that every oracle that computes a probabilistic inversion of $f$ computes $\textup{{0}}^{\prime}$. The converse is:
@@
Each total computable $f:2^{\omega}\to 2^{\omega}$ with positive range has a $\textup{{0}}^{\prime}$-computable positive inversion.
@@
$\forall y\in f(2^{\omega}):\ \Phi(y^{\prime})\in f^{-1}(y).$
By [8, Theorem 8.14.5] we have $y^{\prime}\leq_{T}y\oplus\emptyset^{\prime}$ for almost every $y$. Since $f(2^{\omega})$ is positive, by countable additivity there exists a Turing functional $\Psi$ with
[…]
@@
$\{\hskip 0.85355pt{y\in f(2^{\omega})}:{g(y)\in f^{-1}(y)}\hskip 0.85355pt\}=\{\hskip 0.85355pt{y}:{f(g(y))=y}\hskip 0.85355pt\}\supseteq L.$
Since $\mu(L)>0$ this is a positive subset of $f(2^{\omega})$. ∎
We can now characterize the complexity of positive and probabilistic inversions of total computable functions $f:2^{\omega}\to 2^{\omega}$ with positive range.
@@
The following are equivalent for $w\in 2^{\omega}$:
@@
where $f:2^{\omega}\to 2^{\omega}$ is assumed to have positive range.
@@
By [2, Theorem 4.5] mentioned above we get (ii)$\to$(iii) while (i)$\to$(ii) is immediate. Theorem 2.9 gives (iii)$\to$(i). ∎
It is not hard to show that $\textup{{0}}^{\prime}$ does not compute a positive inversion of some partial computable $f:2^{\omega}\to 2^{\omega}$ with positive range. We show that under a mild condition on $f$ oracle $\textup{{0}}^{\prime\prime}$ is sufficient for this purpose. The condition is that $f$ maps to random reals with positive probability, which is weaker than randomness-preservation and present in Levin’s oneway functions [2].
@@
$\forall y\in f(2^{\omega}):\ \Phi((y\oplus\emptyset^{\prime})^{\prime})\in f^{-1}(y).$
@@
Since $f(2^{\omega})\in\Pi^{0}_{1}(\emptyset^{\prime})$ we have $\mu(f(2^{\omega}))>0$ and by countable additivity there exists a Turing functional $\Psi$ with $\mu(L)>0$ where
$L:=\{\hskip 0.85355pt{y\in f(2^{\omega})}:{\Psi(y\oplus\emptyset^{\prime\prime})=(y\oplus\emptyset^{\prime})^{\prime}}\hskip 0.85355pt\}.$
@@
$\{\hskip 0.85355pt{y\in f(2^{\omega})}:{g(y)\in f^{-1}(y)}\hskip 0.85355pt\}=\{\hskip 0.85355pt{y}:{f(g(y))=y}\hskip 0.85355pt\}\supseteq L.$
Since $\mu(L)>0$ this is a positive subset of $f(2^{\omega})$. ∎

### v3

@@ 1 Introduction
    deterministic and probabilistic computation
@@
    null or positive (measure) domain and range.
@@
Under randomness-preservation (in the sense of Martin-Löf [20]) injective functions are easier to invert. The existence of an injective partial computable oneway function is unknown. A partial answer is given in §4 where a left-oneway (the notion of Gács [11]) partial computable injective function is constructed.
Notation. We use $x,y,\cdots$ to denote reals and $\sigma,\tau,\cdots$ to denote strings. The bit of $x$ at position $n$ is $x(n)$, starting with $n=0$. Let $|\sigma|$ be the length of string $\sigma$. Let $\omega$ be the set of natural numbers, $\{0,1\}^{\omega}$ be the set of all reals, $\{0,1\}^{<\omega}$ the set of finite binary strings and $\{0,1\}^{n}$ the set of binary with length $n$. Let $\preceq$ be the prefix relation between strings and reals, and $\prec$ the strict prefix relation. Let $0^{n}$ be the $n$-bit string of 0s and $0^{\omega}:=000\dots$. For $x,y\in\{0,1\}^{\omega}$ let
  • $x\upharpoonright_{n}$ be the $n$-bit prefix of $x$, so $x\upharpoonright_{n}=x(0)x(1)\cdots x(n-1)$,
  • $x\oplus y$ be the real $z$ with $z(2n)=x(n)$ and $z(2n+1)=y(n)$.
A tree $T$ is a non-empty subset of $\{0,1\}^{<\omega}$ such that if $\sigma\prec\tau\in T$ then $\sigma\in T$. A real $x$ is a path of $T$ if $\forall n,\ x\upharpoonright_{n}\in T$ . Let $[T]$ be the class of all paths of $T$.
The Cantor space is $\{0,1\}^{\omega}$ with the topology generated by basic open sets
$\llbracket\sigma\rrbracket=\{x\in\{0,1\}^{\omega}:\sigma\prec x\}$
for $\sigma\in\{0,1\}^{<\omega}$. The unions of basic open sets called are called open subsets of $\{0,1\}^{\omega}$ and their intersections are called are closed subsets. Let $\mu$ be the uniform measure on $\{0,1\}^{\omega}$ defined by $\mu(\llbracket\sigma\rrbracket)=2^{-|\sigma|}$. We identify $\{0,1\}^{\omega}$ with $\{0,1\}^{\omega}\times\{0,1\}^{\omega}$ via $(x,y)\mapsto x\oplus y$. Classes of reals with positive $\mu$-measure are positive and those with $\mu$-measure 0 are null.
@@
We use the notion of randomness and relative randomness in the sense of Martin-Löf [20] and denote the class of all random reals by $\mathsf{ML}$. We often refer to van Lambalgen’s Theorem [26] (see [8, Corollary 6.9.3]): for every $x,y,z$ the real $x\oplus y$ is $z$-random if and only if $x$ is $z$-random and $y$ is $x\oplus z$-random.
@@ 2 Complexity bounds and inversions
We give complexity bounds on the domain, range and inversions of partial computable functions, assuming familiarity with the arithmetical, hyperarithmetical and analytic hierarchies on $\omega,\{0,1\}^{\omega}$. We write $\Pi^{0}_{1}(w)$ for $\Pi^{0}_{1}$ relative to oracle $w$ and similarly for the other definability classes.
@@ 2.1 Partial computable functions on the reals
A Turing machine with an infinite oracle tape and an infinite one-way output tape can be regarded as an effective map $f$ on the reals: given $x\in\{0,1\}^{\omega}$ on its oracle tape the machine runs indefinitely and if the cells of the output tape contain the digits of $y$ we write $y=f(x)$. This is the standard concept of computability on the reals from computable analysis [23, 27].
Let $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ denote a function from a nonempty subset of $\{0,1\}^{\omega}$ to $\{0,1\}^{\omega}$. Let $\mathsf{dom}(f)$ be the domain and $f(\{0,1\}^{\omega})$ be the range of $f$.
@@
A representation is a partial computable $\hat{f}:\subseteq\{0,1\}^{<\omega}\to\{0,1\}^{<\omega}$ with
@@
and is a representation of $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ if for all $x$ and
$\forall n,\ \hat{f}(x\upharpoonright_{n})\prec f(x)\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf(x)\downarrow\iff\lim_{n}|\hat{f}(x\upharpoonright_{n})|=\infty.$
A function $f$ is $w$-computable if it has a $w$-computable representation.
@@ 2.2 Complexity of domain and range
If $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ is partial computable:
@@
$P=\bigcap_{i}\bigcup_{j}V_{i,j}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptV_{0,0}=\{0,1\}^{\omega},V_{i,j}\subseteq V_{i,j+1}.$
@@
Continuous functions $f$ on subsets of $\{0,1\}^{\omega}$ map compact sets to compact sets. Since $\{0,1\}^{\omega}$ is a Hausdorff space, the same holds for closed sets. An effective version of this fact is true when $f$ is partial computable.
@@
Let $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ be partial computable.
Then for any $w\in\{0,1\}^{\omega}$:
@@
For (i) let $\hat{P}$ be a $w$-computable tree with $\hat{P}]=P$. We show
$y\in f(P)\iff\forall n\ \exists\sigma\in\hat{P}\cap\{0,1\}^{n},\ \hat{f}(\sigma)\prec y.$ (1)
If $y\in f(P)$ then $\exists x\in P,\ f(x)=y$ so $\forall n\ \big{(}{x\upharpoonright_{n}\in\hat{P}\cap\{0,1\}^{n}\ \wedge\ \hat{f}(x\upharpoonright_{n})\prec y}\big{)}$.
[…]
  • since $\sigma_{i}\in\hat{P}$ we get $\tau_{i}\in\hat{P}$ so $x\in[\hat{P}]=P\subseteq\mathsf{dom}(f)$ and $f(x)\downarrow$,
@@
Although the domain of any partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ is arithmetically definable (indeed $\Pi^{0}_{2}$) as a subset of $\{0,1\}^{\omega}$ it is possible that it is nonempty and has no arithmetically definable member.
@@
    if $x\in\Delta^{1}_{1}$ there is partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with domain $\mathsf{dom}(f)=\{z\}$ and $x<_{T}z$
@@
    there is partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with uncountable $\mathsf{dom}(f)$ and $\mathsf{dom}(f)\cap\Delta^{1}_{1}=\emptyset$.
@@
If $x\in\Delta^{1}_{1}$ there is an ordinal $\alpha<\omega_{1}^{\rm CK}$ with $x<_{T}\emptyset^{(\alpha)}$. Since $\{\emptyset^{(\alpha)}\}\in\Pi^{0}_{2}$ by [7, Theorem 2.1.4], by Proposition 2.1 (iii) we get (i). Let
@@
be a computable injective function from $\omega^{\omega}$ into $\{0,1\}^{\omega}$. By [7, Proposition 2.5.2] there is a computable tree $T\subseteq\omega^{\omega}$ such that $[T]$ is uncountable and $T]\cap\Delta^{1}_{1}=\emptyset$. Every $x\in\pi([T])$ computes a member of $[T]$ so $\pi([T])\cap\Delta^{1}_{1}=\emptyset$.
Since $\pi([T])\in\Pi^{0}_{2}$ we get (ii) by Proposition 2.1 (iii). ∎
So a partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ may have a non-empty domain such that all arithmetically definable closed subsets (represented as the set of infinite paths through an arithmetically definable tree) of it are empty.
By Proposition 2.2 (i) the range of any total computable $f:\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ is $\Pi^{0}_{1}$. However, the range of a partial computable $f$ is only $\Sigma^{1}_{1}$. Recall ([7, Definition 1.1.6]) that a class $Q\subseteq\{0,1\}^{\omega}$ is $\Sigma^{1}_{1}$ if for each $y$:
$y\in Q\iff\exists\alpha\in\omega^{\omega}\ \forall s,\ P(\alpha,y,s)$
where $P\subseteq\omega^{\omega}\times\{0,1\}^{\omega}\times\omega$ is a total computable predicate.
Proposition 2.4 ([29]).
The following hold:
  1. (i)
    the range of any partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ is $\Sigma^{1}_{1}$
  2. (ii)
    each $\Sigma^{1}_{1}$ class is the range of a partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$.
Proof..
Clause (i) is straightforward. For (ii), we fix an effective injection from $\omega^{\omega}$ to $\{0,1\}^{\omega}$, for example $\pi(\alpha):=0^{1+\alpha(0)}10^{1+\alpha(1)}\dots 0^{1+\alpha(n)}\dots$.
For any $Q\in\Sigma^{1}_{1}$, let $P\subseteq\omega^{\omega}\times\{0,1\}^{\omega}\times\omega$ be a total computable predicate with
$y\in Q\iff\exists\alpha\in\omega^{\omega}\ \forall s,\ P(\alpha,y,s).$
Let $f$ be such that
$f(x\oplus y)=\begin{cases}y&x=\pi(\alpha)\text{ for some }\alpha\ \land\ \forall s,\ P(\alpha,y,s)\\ \uparrow&\text{otherwise}\end{cases}$
Then $f$ is partial computable and $f(\{0,1\}^{\omega})=Q$. ∎
@@
Let $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ be a partial computable function.
If $f$ has positive domain then:
@@ 2.3 Complexity of inversion
For $f,g:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ we say that $g$ is an inversion of $f$ if
$\forall y\in f(\{0,1\}^{\omega}),\ f(g(y))=y.$
By [2, Corollary 3.2] every total computable injective $f$ has a total computable inversion. We are particularly interested in functions that preserve randomness.
@@
We say that $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ is random-preserving if it has positive domain and $f(x)$ is random for each random $x\in\mathsf{dom}(f)$.
@@
$p^{z}(\langle n,s\rangle)\$:={ $:=\ \begin{cases}\ \ 2n&\textrm{if $n\in z^{\prime}_{s+1}-z^{\prime}_{s}$}\\[5.69046pt] 2\langle n,s\rangle+1&\text{otherwise}\end{cases}$
$h^{z}(x)(\langle n,s\rangle)\$$:=\ x(p^{z}(\langle n,s\rangle))$
@@
is a total computable random-preserving surjection. For contradiction, assume that $g$ is a continuous inversion of $f$. Since $f$ is total, $g$ has $G_{\delta}$ domain and therefore must be $w$-computable for some $w$. Now
@@
Given $f,g:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ we say that $g$ is a
@@
By [2, Theorem 4.5] there is a total computable random-preserving surjection $f$ such that every oracle that computes a probabilistic inversion of $f$ computes $\textup{{0}}^{\prime}$. We show that this complexity bound is optimal.
@@
Each total computable $f:\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with positive range has a $\textup{{0}}^{\prime}$-computable positive inversion.
@@
$\forall y\in f(\{0,1\}^{\omega}):\ \Phi(y^{\prime})\in f^{-1}(y).$
By [8, Theorem 8.14.5] we have $y^{\prime}\leq_{T}y\oplus\emptyset^{\prime}$ for almost every $y$. Since $f(\{0,1\}^{\omega})$ is positive, by countable additivity there exists a Turing functional $\Psi$ with
$\mu(L)>0\hskip 5.69046pt\textrm{where}\hskip 8.5359ptL:=\{\hskip 0.85355pt{y\in f(\{0,1\}^{\omega})}:{\Psi(y\oplus\emptyset^{\prime})=y^{\prime}}\hskip 0.85355pt\}.$
@@
$\{\hskip 0.85355pt{y\in f(\{0,1\}^{\omega})}:{g(y)\in f^{-1}(y)}\hskip 0.85355pt\}=\{\hskip 0.85355pt{y}:{f(g(y))=y}\hskip 0.85355pt\}\supseteq L.$
Since $\mu(L)>0$ this is a positive subset of $f(\{0,1\}^{\omega})$. ∎
We can now characterize the complexity of positive and probabilistic inversions of total computable functions $f:\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with positive range.
@@
The following are equivalent for $w\in\{0,1\}^{\omega}$:
@@
where $f:\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ is assumed to have positive range.
@@
[…]
It is not hard to show that $\textup{{0}}^{\prime}$ does not compute a positive inversion of some partial computable $f:\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with positive range. We show that under a mild condition on $f$ oracle $\textup{{0}}^{\prime\prime}$ is sufficient for this purpose. The condition is that $f$ maps to random reals with positive probability, which is weaker than randomness-preservation and present in Levin’s oneway functions [2].
@@
$\forall y\in f(\{0,1\}^{\omega}):\ \Phi((y\oplus\emptyset^{\prime})^{\prime})\in f^{-1}(y).$
@@
Since $f(\{0,1\}^{\omega})\in\Pi^{0}_{1}(\emptyset^{\prime})$ we have $\mu(f(\{0,1\}^{\omega}))>0$ and by countable additivity there exists a Turing functional $\Psi$ with $\mu(L)>0$ where
$L:=\{\hskip 0.85355pt{y\in f(\{0,1\}^{\omega})}:{\Psi(y\oplus\emptyset^{\prime\prime})=(y\oplus\emptyset^{\prime})^{\prime}}\hskip 0.85355pt\}.$
@@
$\{\hskip 0.85355pt{y\in f(\{0,1\}^{\omega})}:{g(y)\in f^{-1}(y)}\hskip 0.85355pt\}=\{\hskip 0.85355pt{y}:{f(g(y))=y}\hskip 0.85355pt\}\supseteq L.$
Since $\mu(L)>0$ this is a positive subset of $f(\{0,1\}^{\omega})$. ∎